OpenStudy (anonymous):

Approximate the zero of the function in the indicated interval to six decimal places. f(x) = x^3 - x - 1 for my x0 is use 1.5 But my problem is that for every approximation i get a totally different number

OpenStudy (anonymous):

f(x)=3.375-1.5-1=.875

OpenStudy (anonymous):

i am going to bet you are supposed to use newton's method

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes sorry i forgot to indicate that part

OpenStudy (anonymous):

first guess is 1.5 then take \[1.5-\frac{f(1.5)}{f'(.15)}\]

OpenStudy (anonymous):

i know how to do it its just that for every approximation i get a different number

OpenStudy (anonymous):

ok a spread sheet is nice, or we can make a formula. or you can just do it step by step

OpenStudy (anonymous):

yeah i did it in spread sheet form well not on excel or anything but on paper

OpenStudy (anonymous):

\[f(1.5)=.875\] \[f'(1.5)=5.75\] \[1.5-\frac{.875}{5.75}=1.347826087\]

OpenStudy (anonymous):

that is \[x_1\]

OpenStudy (anonymous):

now \[1.347826087-\frac{f(1.347826087)}{f'(1.347826067)}=1.3252\]

OpenStudy (anonymous):

rounded but next two decimals are 0

OpenStudy (anonymous):

that is \[x_2\]

OpenStudy (anonymous):

then you can do it again. alternatively you can just get a formula for \[x-\frac{x^3-x-1}{3x^2-1}=\frac{2x^3+1}{3x^2-1}\]

OpenStudy (anonymous):

and keep plugging the numbers is that

OpenStudy (anonymous):

a pain no matter which way you do it, but i will do the last one for you

OpenStudy (anonymous):

ok u skippedd a step what after i found f(x1) and f'(x1) then i divided the two using the formula f(x1)/f'(x1)

OpenStudy (anonymous):

i mean f(x1)/-f'(x1)

OpenStudy (anonymous):

right and then subtract that from \[x_1\]

OpenStudy (anonymous):

it is \[x_2=x_1-\frac{f(x_1)}{f'(x_1)}\]

OpenStudy (anonymous):

then i do 1.5 minus that

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but we dont get the same answer

OpenStudy (anonymous):

i wrote the answer above. then finally the next one is 1.324718147

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

start with 1.5

OpenStudy (anonymous):

got 1.652173913 for my first approximation

OpenStudy (anonymous):

then subtract \[\frac{f(1.5)}{f'(1.5)}\]

OpenStudy (anonymous):

oh i see what you are doing.

OpenStudy (anonymous):

it is \[xn=x_{n-1}-\frac{f(x_{n-1})}{f'(x_{n-1})}\]

OpenStudy (anonymous):

you had a -f'(x) in the denominator so when you subtracted you really added!

OpenStudy (anonymous):

use the formula i wrote \[x-\frac{f(x)}{f'(x)}\]

OpenStudy (anonymous):

compute f(x), and f'(x) then divide and then subtract

OpenStudy (anonymous):

denominator should just be f'(x) and then subtract. not -f'(x)

OpenStudy (anonymous):

i wrote the answer above and checked it. it gives 0 to 6 decimal places

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

good! you were putting -f'(x) ind the denominator and then subtracting, which is like adding.

OpenStudy (anonymous):

lol my lecturer only did one example so i wasnt sure if that was suppose to be there THANK YOU SATELLITE

OpenStudy (anonymous):

yw!